# Reference Grounded Skill Discovery

### v1

@@ Abstract
Scaling unsupervised skill discovery algorithms to high-DoF agents remains challenging. As dimensionality increases, the exploration space grows exponentially, while the manifold of meaningful skills remains limited. Therefore, semantic meaningfulness becomes essential to effectively guide exploration in high-dimensional spaces. In this work, we present Reference-Grounded Skill Discovery (RGSD), a novel algorithm that grounds skill discovery in a semantically meaningful latent space using reference data. RGSD first performs contrastive pretraining to embed motions on a unit hypersphere, clustering each reference trajectory into a distinct direction. This grounding enables skill discovery to simultaneously involve both imitation of reference behaviors and the discovery of semantically related diverse behaviors. On a simulated SMPL humanoid with $359$-D observations and $69$-D actions, RGSD learns structured skills including walking, running, punching, and side stepping, and also discovers related novel behaviors. In downstream control tasks, RGSD outperforms imitation-based skill acquisition baselines. Our results suggest that lightweight reference-guided grounding offers a practical path to discovering semantically rich and structured skills in high-DoF systems.
@@ 1 Introduction
The ultimate goal of unsupervised skill discovery is to acquire a set of reusable skills that can be applied to arbitrary downstream tasks. Achieving this goal requires the learned skills to satisfy two key desiderata: diversity and semantic meaningfulness. First, the skill set should be sufficiently diverse to cover the wide distribution of possible downstream tasks. For example, a manipulator should be equipped with both push and grasp skills; if it only learns to push, it will fail on many other tasks, such as cup holding or object stacking. Second, the skills must be semantically meaningful, since downstream tasks of interest are typically defined in semantic terms. For instance, a humanoid might acquire a skill that produces a meaningless jittering behavior, such as vibrating both arms in the air. While this behavior may be diverse, it is difficult to imagine a task where such a skill would be useful. Therefore, it is crucial that discovered skills are not only diverse but also semantically interpretable.
@@
Recent progress in skill discovery has demonstrated the ability to learn skills that satisfy both desiderata, particularly in relatively low degree-of-freedom (DoF) systems such as benchmark environments (Eysenbach et al., 2018; Gregor et al., 2016; Laskin et al., 2022), quadrupeds (Cheng et al., 2024; Atanassov et al., 2024; Rho et al., 2025; Cathomen et al., 2025), or simple robotic manipulators with grippers (Park et al., 2023a; Rho et al., ). However, to the best of our knowledge, scaling unsupervised skill discovery to high-DoF agents remains an open challenge. As the DoF increases, the exploration space grows exponentially, while the portion of the semantically meaningful manifold remains relatively small. This mismatch renders skill discovery in high-DoF systems particularly difficult. As shown in Fig. 1, skills learned by the state-of-the-art unsupervised skill discovery algorithm, METRA (Park et al., 2023b), fail to yield meaningful behaviors. Joints move randomly, producing highly unstructured motions in which arms, legs, torso, and head move independently and arbitrarily. This highlights the necessity of a mechanism that grounds skill discovery within a meaningful manifold.
@@
Our key insight is that to tame the curse of dimensionality in high-DoF skill discovery, we need to construct a semantically meaningful skill latent space a priori and constrain exploration within this space. In this paper, we propose Reference-Grounded Skill Discovery (RGSD), a simple yet effective algorithm that achieves this by leveraging reference data to ground the latent skill space before exploration begins. Building upon DIAYN (Eysenbach et al., 2018), RGSD first uses contrastive learning on reference trajectories to embed motions on a unit hypersphere, where each reference behavior corresponds to a distinct direction. This grounding step establishes a semantically meaningful manifold that constrains subsequent exploration. RGSD then utilizes this pre-structured latent space to simultaneously imitate reference skills and discover novel behaviors that are semantically related to the references. This two-stage approach is analogous to recent large language model training regimes (Guo et al., 2025; Yang et al., 2025), where pretraining with self-supervised objectives establishes a meaningful exploration space before reinforcement learning fine-tuning. As a result, RGSD not only reproduces motions present in the reference data but also discovers new skills that emerge as coherent variations of existing ones, rather than arbitrary high-dimensional movements.
@@ 2.1 Unsupervised skill discovery
The most widely adopted approach for acquiring diverse skills is based on maximizing the mutual information (MI) $\mathcal{I}(S;Z)$ between a latent variable $z$ and the agent’s states $s$ (Gregor et al., 2016; Eysenbach et al., 2018; Hansen et al., 2019; Liu & Abbeel, 2021a; Laskin et al., 2022). However, because MI is typically estimated via KL divergence, it suffers from a key limitation: even minor differences between skills can fully maximize the objective, as long as the discriminator can distinguish them. To address this, methods based on the Wasserstein Dependency Measure (WDM, Ozair et al. (2019)) have been proposed (He et al., 2022; Park et al., 2021; 2023b), which explicitly seek maximal differences between skills. Nevertheless, our method builds on MI-based approaches, and in Section 5.4 we explain why extending WDM-based algorithms with our idea faces inherent challenges. Another relevant line of work is CIC (Laskin et al., 2022), which, like our approach, employs contrastive learning. The key difference is that CIC optimizes a contrastive objective on online data gathered by the agent, leaving it with the same scaling difficulties in high-DoF systems.
More recently, approaches such as LGSD (Rho et al., ) and DoDont (Kim et al., 2024) have explored injecting guidance from LLMs or videos to semantically ground skill discovery. While motivated by a similar goal, these methods still struggle to scale to high-DoF systems, as their guidance is provided only at a high level. In contrast, our method leverages reference data directly, enabling a more effective grounding mechanism that scales to complex, high-DoF agents.
@@ 2.2 Imitation-based skill acquisition
Imitation learning represents another major paradigm for acquiring useful set of skills. Before drawing connections to RGSD, we emphasize a key distinction: imitation learning objectives focus on matching the occupancy measure between demonstration data and the learned policy, while RGSD being a discovery algorithm which explicitly encourages visiting states not present in the dataset. This fundamental difference allows RGSD to discover a wider range of skills, as we show empirically in our experiments.
Most imitation learning approaches build on GAIL(Ho & Ermon, 2016), which employs adversarial learning to align the marginal state distribution of the learned policy with that of the reference data. InfoGAIL(Li et al., 2017) extends this framework by introducing an MI loss that improves both controllability and dataset coverage, resulting in a latent-conditioned policy. Based on InfoGAIL, ASE(Peng et al., 2022) successfully scaled skill learning to humanoids robots in simulator by replacing action labels with state transitions $(s,s^{\prime})$ and introducing additional engineering techniques to improve motion quality. CALM(Tessler et al., 2023) further extends ASE by a incorporating a motion encoder and conditional discriminator, enabling the generation of skills conditioned on specific motions and thereby improving fidelity and controllability.
Finally, within the umbrella of zero-shot goal-conditioned RL, Meta-Motivo (Tirinzoni et al., 2025) shares a similar motivation with our work. It employs forward–backward representation learning on reference data to acquire a skill set specialized for zero-shot RL. The key difference is that, unlike RGSD, it does not explicitly aim to discover new skills, which distinguishes our contribution.
@@ 3 Preliminaries
A common formulation is to maximize the mutual information (MI) $\mathcal{I}(S;Z)$ between skills and visited states. Prior work (Gregor et al., 2016; Eysenbach et al., 2018; Laskin et al., 2022; Liu & Abbeel, 2021b) leverages the decomposition $\mathcal{I}(s;z)=\mathcal{H}(s)-\mathcal{H}(s\mid z)=\mathcal{H}(z)-\mathcal{H}(z\mid s).$ Because $\mathcal{H}(z\mid s)$ is intractable, Eysenbach et al. (2018) introduced a neural encoder $q_{\phi}(z|s)$ to obtain a variational lower bound $\mathcal{G}(\theta,\phi)$ of the skill discovery objective $\mathcal{F}(\theta)$:
@@ 4.1 Pretraining: grounding latent space on reference motions
and optimize the InfoNCE loss (Oord et al., 2018):
@@
Here, $\mathrm{sim}(z_{i},z_{j})=z_{i}^{\top}z_{j}$ denotes cosine similarity since all vectors lie on the unit hypersphere, and $T=1/\kappa$ is the temperature. A detailed derivation is provided in Appendix A.
As depicted in Fig.2-pretraining, the mapping of motions in latent space is spread out before training. However when pretraining is completed, we end up with a within-motion alignment, meaning that the mapping of every state $s\in m$ in each motion $m$ points in exactly the same direction. This characteristic is important for the following section; therefore, for completeness, we provide a proof of this claim in Appendix B.
@@ 4.2 Imitation of reference skills
This formulation can be viewed as feature-based imitation with structured representations, as proposed by Li et al. (2025). Unlike DeepMimic (Peng et al., 2018) or MaskedMimic (Tessler et al., 2024), which reward joint-level similarity, our method evaluates similarity in a learned latent space.
@@
For the proposed reward to serve as a valid imitation objective, it must satisfy two key conditions: (1) it should attain its optimum when the agent visits the exact states of the reference motion, and (2) the reward landscape around the reference states should be locally quasi-concave, ensuring that deviations from the reference lead to a monotonic decrease in the reward within a neighborhood of the optimum. We claim that, under the assumption of “perfect alignment within motion” introduced in the previous section, our reward formulation in Eq. 7 satisfies both conditions, and we provide the proof in Appendix C.1. To exploit this local concavity in practice, we apply early termination: whenever the agent deviates from the reference motion beyond a specified threshold measured by cartesian error, the episode is terminated. This mechanism guarantees that the reward functions as a legitimate imitation objective.
@@ 4.3 Discovery of new skills
where $m$ is the motion sampled by RSI. The full algorithmic details are provided in Appendix D.
@@ 5 Experimental results and discussion
Figure 3: Example skills. RGSD generates diverse behaviors when conditioned on different latent vectors. The figure shows motions from a single policy conditioned on distinct latent vectors sampled near the embedding of the (a) running, (b) backward, (b) sidestepping, and (b) punching motion.
In this section, we investigate four questions: (1) Can RGSD imitate reference motions with high fidelity? (2) Can RGSD discover novel skills that are still semantically related to the references? (3) Can the learned skills be effectively leveraged for downstream tasks? (4) Can RGSD be integrated with METRA?
@@
We compare RGSD against both unsupervised skill discovery (USD) methods and imitation learning (IL) based skill acquisition methods. To assess the impact of reference guidance, we include pure USD baselines. Specifically, we compare against DIAYN (Eysenbach et al., 2018) and METRA (Park et al., 2023b). DIAYN, which forms the backbone of RGSD, also serves as an ablation without reference grounding, while METRA represents a distance-maximization approach to USD that explicitly seeks maximal differences between skills. For IL baselines, we consider ASE and CALM. ASE builds on InfoGAIL (Li et al., 2017), which combines mutual information maximization with an adversarial imitation reward, making it the most direct baseline that integrates discovery with imitation. CALM, a variant of ASE, further augments the imitation objective with a motion-conditioned discriminator. We emphasize that both ASE and CALM rely on GAIL-style rewards, whereas our approach introduces a novel imitation reward derived from the DIAYN objective. To focus purely on algorithmic comparison, we excluded the additional engineering techniques applied in the original ASE and CALM implementations. The specific modifications are detailed in Appendix E.2.
@@
We selected 20 reference motions from the ACCAD dataset. These motions are categorized into five tasks: walk, run, sidestep, backward, and punch, with each task containing 2 to 6 relevant motions. A complete list of motions is provided in Appendix E.3. Training was conducted in the GPU-based simulator Isaac Gym (Makoviychuk et al., 2021), using PPO (Schulman et al., 2017) as the RL algorithm. The full set of hyperparameters is provided in Appendix E.4. To ensure stable training, all methods employed an early termination condition: whenever the robot fell, the episode was terminated.
@@ 5.1 Evaluation of Imitation
For CALM, selecting the right latent that represents each motion is straightforward since it also includes motion encoders. For methods without encoders, we uniformly sample 500 latent vectors, select the one that minimizes Cartesian error, and re-evaluate using this vector to ensure fairness. We found that 500 samples were sufficient, as increasing the number further did not yield noticeable improvements.
@@
As shown in Table 1, RGSD achieves both low Cartesian error and low FID scores across most tasks. In contrast, pure USD baselines fail to discover skills that closely resemble the reference dataset. While DIAYN has demonstrated the ability to discover basic locomotion behaviors in low DoF ranging from 3 to 6 environments such as HalfCheetah, Ant, and Hopper (Brockman et al., 2016), they struggle with the 69-DoF SMPL agent, producing behaviors that are not semantically meaningful. This emphasizes the impact of reference guidance.
When compared to CALM, the results highlight a trade-off between fidelity and naturalness. CALM achieves lower FID scores in 3 out of 5 tasks, indicating that its motions generally appear more natural. However, RGSD outperforms CALM in Cartesian error on 4 out of 5 tasks, demonstrating higher trajectory fidelity. This contrast reflects a key difference in objective design: CALM employs an occupancy-matching objective at the motion level, leading to smoother but less precise reproductions, whereas RGSD relies on frame-level similarity rewards, yielding high fidelity motions.

### v2

@@ Abstract
Scaling unsupervised skill discovery algorithms to high-DoF agents remains challenging. As dimensionality increases, the exploration space grows exponentially, while the manifold of meaningful skills remains limited. Therefore, semantic meaningfulness becomes essential to effectively guide exploration in high-dimensional spaces. In this work, we present Reference-Grounded Skill Discovery (RGSD), a novel algorithm that grounds skill discovery in a semantically meaningful latent space using reference data. RGSD first performs contrastive pretraining to embed motions on a unit hypersphere, clustering each reference trajectory into a distinct direction. This grounding enables skill discovery to simultaneously involve both imitation of reference behaviors and the discovery of semantically related diverse behaviors. On a simulated SMPL humanoid with $359$-D observations and $69$-D actions, RGSD successfully imitates skills such as walking, running, punching, and sidestepping, while also discover variations of these behaviors. In downstream locomotion tasks, RGSD leverages the discovered skills to faithfully satisfy user-specified style commands and outperforms imitation-learning baselines, which often fail to maintain the commanded style. Overall, our results suggest that lightweight reference-grounding offers a practical path to discovering semantically rich and structured skills in high-DoF systems.
@@ 1 Introduction
The ultimate goal of unsupervised skill discovery is to acquire a set of reusable skills that can be applied to arbitrary downstream tasks. Achieving this goal requires the learned skills to satisfy two key desiderata: diversity and semantic meaningfulness. First, the skill set should be sufficiently diverse to cover the wide distribution of possible downstream tasks. For example, a humanoid agent should be equipped with both turn-left and turn-right skills; if it only learns to walk-forward, it will fail on general locomotion tasks, such as goal reaching or heading. Second, the skills must be semantically meaningful, since downstream tasks of interest are typically defined in semantic terms. For instance, a humanoid might acquire a skill that produces a meaningless jittering behavior, such as vibrating both arms in the air. While this behavior may be diverse, it is difficult to imagine a task where such a skill would be useful. Therefore, it is crucial that discovered skills are not only diverse but also semantically interpretable.
@@
Recent progress in skill discovery has demonstrated the ability to learn skills that satisfy both desiderata, particularly in relatively low degree-of-freedom (DoF) systems such as benchmark environments (diayn; vic; cic), quadrupeds (cheng2024learning; atanassov2024constrained; rho2025unsupervised; cathomen2025divide), or simple robotic manipulators with grippers (park2023controllability; rholanguage). However, to the best of our knowledge, scaling unsupervised skill discovery to high-DoF agents remains an open challenge. As the DoF increases, the exploration space grows exponentially, while the portion of the semantically meaningful manifold remains relatively small. This mismatch renders skill discovery in high-DoF systems particularly difficult. As shown in Fig. 1, skills learned by the state-of-the-art unsupervised skill discovery algorithm, METRA (metra), fail to yield meaningful behaviors. Joints move randomly, producing highly unstructured motions in which arms, legs, torso, and head move independently and arbitrarily. This highlights the necessity of a mechanism that grounds skill discovery within a meaningful manifold.
@@
Our key insight is that to tame the curse of dimensionality in high-DoF skill discovery, we need to construct a semantically meaningful skill latent space a priori and constrain exploration within this space. In this paper, we propose Reference-Grounded Skill Discovery (RGSD), a simple yet effective algorithm that achieves this by leveraging reference data to ground the latent skill space before exploration begins. Building upon DIAYN (diayn), RGSD first uses contrastive learning on reference trajectories to embed motions on a unit hypersphere, where each reference behavior corresponds to a distinct direction. This grounding step establishes a semantically meaningful manifold that constrains subsequent exploration. RGSD then utilizes this pre-structured latent space to simultaneously imitate reference skills and discover novel behaviors that are semantically related to the references. This two-stage approach is analogous to recent large language model training regimes (guo2025deepseek; yang2025qwen3), where pretraining with self-supervised objectives establishes a meaningful exploration space before reinforcement learning fine-tuning. As a result, RGSD not only reproduces motions present in the reference data but also discovers new skills that emerge as coherent variations of existing ones, rather than arbitrary high-dimensional movements.
@@ 2.1 Unsupervised skill discovery
The most widely adopted approach for acquiring diverse skills is based on maximizing the mutual information (MI) $\mathcal{I}(S;Z)$ between a latent variable $z$ and the agent’s states $s$ (vic; diayn; hansen2019fast; liu2021aps; cic). However, because MI is typically estimated via KL divergence, it suffers from a key limitation: even minor differences between skills can fully maximize the objective, as long as the discriminator can distinguish them. To address this, methods based on the Wasserstein Dependency Measure (WDM, ozair2019wasserstein) have been proposed (he2022wasserstein; lsd; metra), which explicitly seek maximal differences between skills. Nevertheless, our method builds on MI-based approaches, and in Section 6 we explain why extending WDM-based algorithms with our idea faces inherent challenges. Another relevant line of work is CIC (cic), which, like our approach, employs contrastive learning. The key difference is that CIC optimizes a contrastive objective on online data gathered by the agent, leaving it with the same scaling difficulties in high-DoF systems.
More recently, approaches such as LGSD (rholanguage) and DoDont (kim2024s) have explored injecting guidance from LLMs or videos to semantically ground skill discovery. While motivated by a similar goal, these methods still struggle to scale to high-DoF systems, as their guidance is provided only at a high level. In contrast, our method leverages reference data directly, enabling a more effective grounding mechanism that scales to complex, high-DoF agents.
@@ 2.2 Imitation-based skill acquisition
Imitation learning offers another major paradigm for acquiring a repertoire of skills. A common approach is based on GAIL (ho2016generative), which uses adversarial training to align the marginal state distribution of the learned policy with that of expert demonstrations. Building on this idea, ASE (peng2022ase) learns a controllable skill embedding from large-scale human motion data using a GAIL-style objective. CALM (tessler2023calm) mitigates the mode collapse observed in ASE by incorporating a conditional discriminator and a motion encoder, enabling skill generation conditioned on specific reference motions. More recently, Meta-Motivo (tirinzoni2025zero) combines GAIL with forward–backward representation learning (blier2021learning; touati2021learning) to extract a skill set tailored for zero-shot RL.
In contrast to these imitation-based approaches, we highlight a key conceptual difference: imitation objectives aim to match the state visitation measure of the demonstrations, whereas RGSD is a discovery algorithm that explicitly encourages visiting states not present from the dataset. This fundamental difference allows RGSD to discover a wider range of skills, as demonstrated in our experiments.
@@ 3 Preliminaries
A common formulation is to maximize the mutual information (MI) $\mathcal{I}(S;Z)$ between skills and visited states. Prior work (vic; diayn; cic; liu2021behavior) leverages the decomposition $\mathcal{I}(s;z)=\mathcal{H}(s)-\mathcal{H}(s\mid z)=\mathcal{H}(z)-\mathcal{H}(z\mid s).$ Because $\mathcal{H}(z\mid s)$ is intractable, diayn introduced a neural encoder $q_{\phi}(z|s)$ to obtain a variational lower bound $\mathcal{G}(\theta,\phi)$ of the skill discovery objective $\mathcal{F}(\theta)$:
@@ 4.1 Pretraining: grounding latent space on reference motions
and optimize the InfoNCE loss (oord2018representation):
@@
Here, $\mathrm{sim}(z_{i},z_{j})=z_{i}^{\top}z_{j}$ denotes cosine similarity since all vectors lie on the unit hypersphere, and $T=1/\kappa$ is the temperature. A detailed derivation is provided in Appendix B.
As depicted in Fig.2-pretraining, the mapping of motions in latent space is spread out before training. However when pretraining is completed, we end up with a within-motion alignment, meaning that the mapping of every state $s\in m$ in each motion $m$ points in exactly the same direction. This characteristic is important for the following section; therefore, for completeness, we provide a proof of this claim in Appendix C.
@@ 4.2 Imitation of reference skills
This formulation can be viewed as feature-based imitation with structured representations, as proposed by li2025feature. Unlike DeepMimic (peng2018deepmimic) or MaskedMimic (tessler2024maskedmimic), which reward joint-level similarity, our method evaluates similarity in a learned latent space.
@@
For the proposed reward to serve as a valid imitation objective, it must satisfy two key conditions: (1) it should attain its optimum when the agent visits the exact states of the reference motion, and (2) the reward landscape around the reference states should be locally quasi-concave, ensuring that deviations from the reference lead to a monotonic decrease in the reward within a neighborhood of the optimum. We claim that, under the assumption of “perfect alignment within motion” introduced in the previous section, our reward formulation in Eq. 7 satisfies both conditions, and we provide the proof in Appendix D.1. To exploit this local concavity in practice, we apply early termination: whenever the agent deviates from the reference motion beyond a specified threshold measured by cartesian error, the episode is terminated. This mechanism guarantees that the reward functions as a legitimate imitation objective.
@@ 4.3 Discovery of new skills
where $m$ is the motion sampled by RSI. The full algorithmic details are provided in Appendix F.
@@ 5 Experimental results and discussion
Figure 3: Example skills. RGSD generates diverse behaviors when conditioned on different latent vectors. The figure shows motions from a single policy conditioned on distinct latent vectors sampled near the embedding of the (a) running, (b) backward, (c) sidestepping, and (d) punching motion.
In this section, we investigate four questions: (1) Can RGSD imitate reference motions with high fidelity? (2) Can RGSD discover novel skills that are still semantically related to the references? (3) Can we control degree of diversity at test time? (4) Can the learned skills be effectively leveraged for downstream tasks?
@@
We compare RGSD against both unsupervised skill discovery (USD) methods and imitation learning (IL) based skill acquisition methods. To assess the impact of reference guidance, we include pure USD baselines. Specifically, we compare against DIAYN (diayn) and METRA (metra). DIAYN, which forms the backbone of RGSD, also serves as an ablation without reference grounding, while METRA represents a distance-maximization approach to USD that explicitly seeks maximal differences between skills.
For the IL baselines, we consider ASE, CALM, and Meta-Motivo, all of which utilize a GAIL-style objective (ho2016generative) for imitation. ASE builds on InfoGAIL (li2017infogail), combining mutual-information maximization with an adversarial imitation reward, making it the closest baseline that jointly addresses skill discovery and imitation. CALM, a variant of ASE, further augments the imitation objective using a motion-conditioned discriminator. Meta-Motivo integrates imitation rewards with forward–backward representation learning.
We emphasize that while these methods rely on GAIL-based imitation rewards, our approach employs a novel imitation reward derived from the DIAYN objective. To ensure a clean algorithmic comparison, we remove additional engineering techniques used in the original ASE and CALM implementations. The exact modifications are described in Appendix G.2.
@@
We selected 20 reference motions from the osu_accad dataset. These motions are categorized into five tasks: walk, run, sidestep, backward, and punch, with each task containing 2 to 6 relevant motions. A complete list of motions is provided in Appendix G.3. Training was conducted in the GPU-based simulator Isaac Gym (makoviychuk2021isaac), using PPO (ppo) as the RL algorithm. The full set of hyperparameters is provided in Appendix G.4. To ensure stable training, all methods employed an early termination condition: whenever the robot fell, the episode was terminated.
@@ 5.1 Evaluation of Imitation
For CALM and Meta-Motivo, selecting the right latent that represents each motion is straightforward since it also includes motion encoders. For methods without encoders, we uniformly sample 500 latent vectors, select the one that minimizes Cartesian error, and re-evaluate using this vector to ensure fairness. We found that 500 samples were sufficient, as increasing the number further did not yield noticeable improvements.
@@
As shown in Table 1, RGSD achieves low Cartesian error while Meta-Motivo achieves low FID scores across most motions. In contrast, pure USD baselines fail to discover skills that closely resemble the reference dataset. While DIAYN has demonstrated the ability to discover basic locomotion behaviors in low DoF ranging from 3 to 6 environments such as HalfCheetah, Ant, and Hopper (brockman2016openai), they struggle with the 69-DoF SMPL agent, producing behaviors that are not semantically meaningful. This emphasizes the impact of reference guidance.
When compared to Meta-Motivo, the results highlight a trade-off between fidelity and naturalness. Meta-Motivo achieves lower FID scores in 4 out of 5 tasks, indicating that its motions generally appear more natural. However, RGSD outperforms Meta-Motivo in Cartesian error on 4 out of 5 tasks, demonstrating higher trajectory fidelity. This contrast reflects a key difference in objective design: Meta-Motivo employs an occupancy-matching objective at the motion level, leading to smoother but less precise reproductions, whereas RGSD relies on frame-level similarity rewards, yielding high fidelity motions.